### v2

# A Systematic Survey on Large Language Models for Evolutionary Optimization: From Modeling to Solving

Abstract
Large Language Models (LLMs), with their strong understanding and reasoning capabilities, are increasingly being explored for tackling optimization problems, especially in synergy with evolutionary computation [^1] This survey primarily focuses on evolutionary optimization, i.e., optimization based on evolutionary computation. For brevity, we use the term optimization throughout to denote this scope.. Despite rapid progress, however, the field still lacks a unified synthesis and a systematic taxonomy. This survey addresses this gap by providing a comprehensive review of recent developments and organizing them within a structured framework. We classify existing research into two main stages: LLMs for optimization modeling and LLMs for optimization solving [^2] In this survey, the terms optimization modeling and optimization solving are used as concise forms of optimization problem modeling and optimization problem solving, respectively.. The latter is further divided into three paradigms according to the role of LLMs in the optimization workflow: LLMs as stand-alone optimizers, low-level LLMs embedded within optimization algorithms, and high-level LLMs for algorithm selection and generation. For each category, we analyze representative methods, distill technical challenges, and examine their interplay with traditional approaches. We also review interdisciplinary applications spanning the natural sciences, engineering, and machine learning. By contrasting LLM-driven and conventional methods, we highlight key limitations and research gaps, and point toward future directions for developing self-evolving agentic ecosystems for optimization. An up-to-date collection of related literature is maintained at https://github.com/ishmael233/LLM4OPT.
I Introduction
Optimization techniques have become indispensable for addressing complex problems across diverse domains, including engineering design [1], economic planning [2], and scientific discovery [3]. In practice, real-world challenges are typically abstracted into mathematical optimization models, which are then solved by optimization algorithms. Broadly, these algorithms can be divided into two categories: exact methods and approximate methods [4]. Exact methods, while theoretically capable of guaranteeing global optimality, often suffer from exponential growth in computational complexity, rendering them impractical for large-scale problems. In contrast, approximate methods, such as heuristics [4], can efficiently tackle large and complex problems, though without convergence guarantees. Moreover, the No Free Lunch (NFL) theorem [5] dictates that no single optimization algorithm can consistently outperform others across all problem classes. Consequently, the practical deployment of optimization algorithms typically demands substantial expertise for configuration [6], selection [7], and even customized design [8]. This heavy reliance on domain knowledge remains a critical barrier to transferring optimization technology from theory to practice.
Against this backdrop, machine learning has emerged as a promising approach to alleviate the complexity of algorithm configuration and reduce dependence on expert knowledge. Among these methods, Reinforcement Learning (RL) [9, 10], with its flexible decision-making framework, has been widely applied at multiple levels of optimization algorithms [11]. At the internal structural level, RL has been employed for population initialization [12], fitness evaluation [13], operator selection [14], and parameter control [15]. At the higher strategic level, RL has enabled automated algorithm selection [16] and even algorithm generation [17]. However, a persistent limitation is that most RL-based approaches are trained on narrowly defined problem sets, which severely restricts their generalization. When confronted with new problems, retraining is often costly and time-consuming. These limitations motivate the exploration of Large Language Models (LLMs) [18], whose powerful semantic understanding, reasoning, and generalization capabilities open new opportunities for optimization research.
The rapid advancements of LLMs are fundamentally reshaping research paradigms across disciplines. Trained on vast corpora, LLMs exhibit strong capabilities in language understanding [19], complex logical reasoning [20], and open-ended knowledge generalization [21]. Their influence extends far beyond Natural Language Processing (NLP) [22], demonstrating transformative potential in high-value domains such as medicine [23] and finance [24]. For example, Huawei’s PanGu Drug Model [25] accelerates drug discovery pipelines, while BloombergGPT [26] integrates diverse financial data to provide domain-specific insights. Within optimization, the semantic and reasoning strengths of LLMs open unique opportunities: on the modeling side, they can translate natural language descriptions into mathematical optimization models [27], thereby advancing the automation and intelligence of the entire optimization workflow; on the solving side, they can function as stand-alone optimizers [28] or collaborate with existing algorithms to enhance performance; .
Figure 1: Overall organization of this survey.
Despite their significant potential, the use of LLMs in optimization is still at an early stage. Existing studies have already revealed important limitations. For example, LLMs as standalone optimizers often fail to surpass classical algorithms in performance [29], and although fine-tuning with domain-specific data can enhance task outcomes, constructing such datasets remains costly and expertise-intensive [27]. Against this backdrop, this survey provides a systematic analysis and guide for this emerging interdisciplinary field. The main contributions are summarized as follows:
  • We present a comprehensive literature review of research on LLMs for optimization, systematically summarizing recent advances, representative methods, and key applications.
  • We propose a structured taxonomy that categorizes existing work into optimization modeling and optimization solving, with the latter further divided into three paradigms: LLMs as optimizers, low-level LLMs embedded in algorithms, and high-level LLMs for algorithm selection and generation. This taxonomy offers a unified framework for understanding and organizing the research landscape.
  • We contrast LLM-driven approaches with traditional methods to identify fundamental challenges and research gaps, and outline promising future directions toward building an intelligent optimization ecosystem.
The overall structure of this survey is illustrated in Fig. 1. Section II reviews related surveys and introduces our taxonomy, followed by Section III, which provides the necessary background on key technologies. Sections IV and V form the core of this work, focusing on LLMs for optimization modeling and solving, respectively. Section VI then examines interdisciplinary applications of these approaches. Section VII outlines future research directions, and Section VIII concludes with a summary of key findings.
TABLE I: Comparison of related surveys on LLMs for optimization. Our work provides the first systematic coverage of all four categories: LLMs for Modeling (LM), LLMs as Optimizers (LO), Low-level LLMs for Optimization Algorithms ($L$), and High-level LLMs for Optimization Algorithms (HL).
Ref. Venue LM LO LL HL Wu et al. [30] TEVC, 2024 Huang et al. [31] SWEVO, 2024 Chao et al. [32] Research, 2024 Yu et al. [33] arXiv, 2024 Liu et al. [34] arXiv, 2024 Ma et al. [35] TEVC, 2024 Our work arXiv, 2025
II Related Surveys and Taxonomy
Several surveys and reviews have examined the use of LLMs in optimization. Table I summarizes their scope and contrasts them with our study. Wu et al. [30] and Huang et al. [31] reviewed the field from a bidirectional perspective, considering both LLM-enhanced optimization algorithms and optimization algorithm-enhanced LLMs. However, in the former direction, their discussions remained limited to scenarios where LLMs serve as solvers or for algorithm generation, without analyzing their broader relationship to the optimization process. Yu et al. [33] adopted a similar classification and proposed a general framework for LLM-assisted optimization. Chao et al. [32] focused specifically on Evolutionary Algorithms (EAs) [36, 37], highlighting component-level analogies between EAs and LLMs to suggest bidirectional knowledge transfer, but their coverage was restricted to cases where LLMs act as operators. Other related works took slightly different perspectives. For instance, Liu et al. [34] provided a broad overview of LLMs in algorithm design, categorizing their roles as optimizers, predictors, extractors, and designers. Ma et al. [35] surveyed meta-black-box optimization, which mentioned LLMs but primarily emphasized conventional reinforcement learning approaches. Overall, existing reviews have concentrated mainly on the solving stage of optimization, with limited attention to the modeling stage. Even in [34], modeling was only briefly introduced as a minor application.
In contrast to existing surveys, this paper provides a systematic review of the entire modeling-to-solving workflow of LLMs for optimization. We propose a new taxonomy that articulates the relationship between LLMs and the optimization process. Specifically, we divide the field into two main categories: LLMs for optimization modeling and LLMs for optimization solving. The latter is further subdivided into three paradigms: LLMs as optimizers, low-level LLM-assisted optimization algorithms, and high-level LLM-assisted optimization algorithms. The definitions of these categories are as follows:
  • LLMs for Optimization Modeling aims to automatically transform unstructured natural language descriptions of problems into mathematical optimization models that can be interpreted and solved by machines. This area represents the foundational step toward a fully LLM-driven optimization workflow, with the central challenge being the transition from ambiguous natural language to precise mathematical formulations. Existing approaches are mainly divided into prompt-based and learning-based methods. Prompt-based approaches typically rely on carefully designed prompting strategies or multi-agent collaboration, offering fast deployment without additional training. In contrast, learning-based approaches often require data synthesis and fine-tuning, which incur training costs but can significantly improve reliability in complex scenarios.
  • LLMs as Optimizers treat LLMs themselves as general-purpose optimizers that solve optimization problems through iterative natural language interaction, independent of traditional algorithmic frameworks. This paradigm leverages the in-context learning and reasoning abilities of LLMs to explore optimization trajectories and generate promising candidate solutions. Although this approach faces challenges such as model dependency and scalability, it represents the earliest and most direct attempt to connect large models with optimization tasks, highlighting its pioneering significance.
  • Low-level LLM-assisted Optimization Algorithms embed LLMs as intelligent components within traditional optimization algorithms to enhance specific operations. Unlike the standalone optimizer paradigm, this approach integrates LLMs tightly with established frameworks, exploiting their analytical and domain knowledge to improve performance. Typical applications include population initialization, operator design, parameter control, configuration, and fitness evaluation. By empowering traditional components to adapt more intelligently to diverse problem characteristics and search states, this paradigm improves both efficiency and solution quality at the algorithmic core.
  • High-level LLM-assisted Optimization Algorithms focus on top-level orchestration rather than internal components. They generally fall into two subcategories: algorithm selection and algorithm generation. Algorithm selection leverages LLMs to choose the most suitable algorithm from a pool for different problem instances, while algorithm generation goes further by enabling LLMs to autonomously design new algorithms tailored to specific tasks. By granting LLMs a global perspective over the entire optimization workflow, this paradigm elevates their role from internal assistant to top-level designer.
III Background
Given that current research on LLMs for internal component control, high-level orchestration, and algorithm generation is predominantly centered on EAs, we first provide the necessary background. Section III-A introduces the framework and core paradigms of EAs, while Section III-B outlines the underlying architectures and key technologies of LLMs.
Figure 2: Technological dependencies of LLMs for optimization, including EA paradigm and workflow, LLM architecture, and related enabling technologies.
III-A Evolutionary Algorithms
EAs are a class of population-based, gradient-free optimization methods inspired by the principles of biological evolution. As illustrated in Fig. 2, the general framework of an EA begins with the initialization of a population of candidate solutions, whose representations can range from numerical vectors to neural networks. Within each iteration, individuals are evaluated using a fitness function. Parent individuals are then selected to produce offspring through genetic operators such as crossover and mutation. A new population is formed by selecting the fittest individuals from both parents and offspring. This evolutionary cycle repeats until a termination criterion is reached, at which point the best-performing individual is returned as the final solution.
Over decades of development, researchers have proposed a diverse set of classical paradigms [38, 39]. Prominent examples include Genetic Algorithms (GAs) [40], Genetic Programming (GP) [41], Differential Evolution (DE) [42], and Particle Swarm Optimization (PSO) [43]. Although these algorithms share the population-based iterative framework, they differ in search mechanisms and application domains. For instance, GAs simulate natural selection and genetic operators, making them well-suited for discrete and combinatorial problems; DE and PSO excel in continuous optimization tasks through distinctive mutation and information-sharing strategies; while GP, with its tree-structured encoding, has been widely applied to automatic programming and symbolic regression. Collectively, these paradigms established the theoretical foundation of evolutionary optimization and provided diverse strategies for tackling increasingly complex problems.
Compared to traditional exact optimization methods, EAs offer significant advantages in handling non-convex, multi-modal, and noisy problems. Their applications span a wide range of domains, including aerospace design [44] and hyperparameter tuning in deep learning [45]. Nonetheless, in line with the NFL theorem, no single EA can achieve universal superiority across all problems at low computational cost. As a result, practitioners often need to devote substantial effort to selecting, configuring, or designing suitable EA variants. To address these challenges, recent research has increasingly integrated EAs with machine learning methods [46]. Notably, reinforcement learning has been combined with evolutionary frameworks, giving rise to approaches such as MOEA/D-DQN [47] and SYMBOL [17]. With the emergence of LLMs, cutting-edge studies are further exploring how LLMs can empower EAs [30], extending from problem formulation [27] to automated algorithm design [34].
III-B Large Language Models
The foundation of LLMs lies in the Transformer architecture [48], which captures global input-output dependencies through self-attention. Building on this architecture, researchers have developed three primary paradigms: encoder-decoder, encoder-only, and decoder-only models [49]. Encoder-decoder models, exemplified by BART [50] and T5 [51], follow the original Transformer design and excel at sequence-to-sequence tasks. Encoder-only models, such as BERT [52], focus on contextual representation learning for semantic understanding. Decoder-only models, including ChatGPT [53] and DeepSeek [54], leverage autoregressive generation for open-domain tasks and currently dominate LLM-for-Optimization research due to their strong generative and reasoning capabilities. Relevant applications mainly employ two pathways: prompt engineering [55] and fine-tuning [56].
Prompt engineering steers LLM outputs by carefully designed instructions without modifying model parameters. Early forms include zero-shot and few-shot prompting [57], which exploit pre-trained knowledge directly or via limited in-context examples. More advanced approaches incorporate structured reasoning, such as Chain-of-Thought (CoT) [20], which enhances performance on complex reasoning tasks by eliciting step-by-step logical inference. CoT has since expanded into richer paradigms like Tree-of-Thought (ToT) [58], Graph-of-Thought (GoT) [59], and domain-specific variants such as Chain-of-Code (CoC) [60] and Chain-of-Knowledge (CoK) [61].
Fine-tuning, in contrast, updates model parameters on task-specific data, enabling deeper adaptation at the cost of labeled data and computational resources. Two strategies are particularly relevant. Instruction fine-tuning [62] reformulates diverse tasks into instruction–input–output triplets, unifying them under an instruction-following paradigm. This significantly improves zero-shot generalization, as demonstrated by Flan-T5 [63]. Alignment fine-tuning ensures model behavior aligns with human intentions and values. Reinforcement Learning with Human Feedback (RLHF) [64] achieves this by optimizing a reward model built from human preferences, while Direct Preference Optimization (DPO) [65] simplifies the process by directly optimizing for preferred outputs. Together, these techniques are central to building safe, reliable, and useful LLMs.
IV LLMs for Optimization Modeling
Optimization modeling is a crucial step in the optimization workflow but traditionally relies heavily on expert knowledge and experience. Recently, LLMs have opened new opportunities for automated optimization modeling. In this section, we introduce two core paradigms: prompt-based methods and learning-based methods. Prompt-based methods (Section IV-A) guide LLMs through carefully designed instructions, often implemented via two-stage workflows, multi-agent collaboration, or interactive prompting. Learning-based methods (Section IV-B), in contrast, fine-tune LLMs to directly generate mathematical formulations, typically requiring large-scale synthetic data for training but offering greater reliability in complex tasks. Finally, we discuss open challenges in this field (Section IV-C). Representative studies are summarized in Table II, while the main paradigms are illustrated in Fig. 3.
Figure 3: Illustration of LLMs for optimization modeling. Approaches can be broadly divided into two categories: (i) prompt-based methods, typically implemented via two-stage prompting, multi-agent collaboration, or interactive frameworks; and (ii) learning-based methods, which generally follow a workflow of data synthesis, model fine-tuning, and evaluation.
TABLE II: Representative works on LLMs for optimization modeling, categorized into prompt-based and learning-based methods.
Method Venue Type Technical Summary Prompt-based Methods Ner4OPT [66] CPAIOR, 2023 Two-stage Fine-tune models for named entity recognition by incorporating traditional NLP methods. AOMG [67] GECCO, 2023 Direct Directly leverage LLMs to generate mathematical models. HG 2.0[68] arXiv, 2023 Two-stage Embed LLMs within the two-stage framework. AMGPT [69] arXiv, 2023 Two-stage Employ fine-tuned models for constraint classification. CoE [70] ICLR, 2023 Multi-agent Construct dynamic reasoning chains using 11 expert agents. OptiMUS [71] ICML, 2023 Multi-agent Develop a conductor agent to coordinate modeling, programming, and evaluation processes. MAMS [72] INFOR, 2024 Multi-agent Develop inter-agent cross-validation to replace solver-dependent verification. EC [73] SGC, 2024 Two-stage Apply the two-stage framework to energy management systems. CAFA [74] NeurIPS, 2024 Two-stage Enhance modeling performance through code-based formalization. MAMO [75] NAACL, 2024 Two-stage Develop the MAMO benchmark with ordinary differential equation extensions. NL2OR [76] arXiv, 2024 Two-stage Predefine abstract structural constraints to regulate the LLM outputs. TRIP-PAL [77] arXiv, 2024 Two-stage Apply the two-stage framework to travel planning. OptLLM [78] arXiv, 2024 Interactive Support both single-input and interactive-input modes. OptiMUS-0.3 [79] arXiv, 2024 Multi-agent Introduce self-corrective prompts and structure-aware modeling based on OptiMUS. MeetMate [80] TiiS, 2024 Interactive Develop an interactive system for user input processing with five selectable task options. VL [81] CUI, 2025 Interactive Convert user priorities into optimization constraints during dialog interactions. LLM-MCTS [82] arXiv, 2025 Two-stage Conduct hierarchical Monte Carlo tree search over the hypothesis space. EquivaMap [83] arXiv, 2025 Two-stage Generate variable mapping functions via LLMs with lightweight verification. MAP [84] arXiv, 2025 Multi-agent Deploy multiple independent reviewers to evaluate modeling results. ORMind [85] arXiv, 2025 Multi-agent Replace the conductor agent with structured, predictable workflows. LEAN-LLM-OPT [86] ssrn, 2025 Multi-agent Integrate RAG to enhance problem classification, and construct instances with query operations. Learning-based Methods LM4OPT [87] INFOR, 2024 Fine-tuning Progressively fine-tune models on the NL4OPT dataset. ORLM [27] OR, 2024 Data synthesis Synthesize data through expansion and augmentation and fine-tune open-source models. ReSocratic [88] ICLR, 2024 Data synthesis Propose inverse data synthesis methodology and construct the OPTIBENCH benchmark. LLMOPT [89] ICLR, 2024 Fine-tuning Introduce model alignment and self-correction mechanisms to mitigate hallucination phenomena. BPP-Search [90] arXiv, 2024 Data synthesis Solve the problem of missing detailed in data synthesis. OptMATH [91] arXiv, 2025 Data synthesis Develop a scalable bidirectional data synthesis approach. LLMBO [92] arXiv, 2025 Data synthesis Propose a method for fine-tuning cost-effective LLMs to tackle specific business challenges. SIRL [93] arXiv, 2025 Fine-tuning Integrate external optimization solvers as verifiable reward validators for reinforcement learning. Step-Opt [94] arXiv, 2025 Data synthesis Increase problem complexity through an iterative problem generation approach. DPLM [95] arXiv, 2025 Data synthesis Combine the diversity of forward generation and the reliability of inverse generation. OptiTrust [96] arXiv, 2025 Data synthesis Develop a verifiable synthetic data generation pipeline.
IV-A Prompt-based Methods
Automated optimization modeling, which aims to formalize real-world challenges into mathematical models, has recently emerged as a significant research direction. To mitigate the reliance on expert experience during this process, researchers have explored various approaches. For instance, Pawlak et al. preliminarily demonstrated the feasibility of automated modeling by analyzing feasible solution samples [97] or employing evolutionary algorithms [98] to generate constraints. However, these methods were often constrained by sample quality and limited in scope [69], primarily focusing on constraint generation rather than a direct conversion from natural language descriptions to complete mathematical models.
To address this research gap, Ramamonjison et al. proposed OptGen [99], the first method to convert unstructured natural language problem descriptions into mathematical optimization models, and subsequently catalyzed the field’s growth with the NL4OPT competition [100]. This competition decomposed the task into two core stages: first, using Named Entity Recognition (NER) to extract key entities such as variables, parameters, and constraints from the problem text, and then generating a complete mathematical model from the entity-annotated description. However, these earlier works largely relied on lightweight pre-trained models like BERT [101, 102] and BART [103, 104]. In contrast, with the advent of LLMs, B. Almonacid et al. were the first to attempt automated modeling using powerful LLMs such as GPT-3.5 [67]. They employed a single-step modeling approach, directly prompting the LLM to output the mathematical model corresponding to a problem description and verifying its correctness with a solver. Nevertheless, while this single-step method proves effective for simple problems, its accuracy and robustness are generally insufficient for meeting the high standards required for complex optimization modeling.
Inspired by the aforementioned studies, researchers have progressively adopted large language models for a two-stage approach involving NER and automated modeling in the NL4OPT task. The Holy Grail 2.0 framework [68] proposed by Tsouros et al., for instance, recommends using specialized models like Ner4OPT for the NER stage. It then further decomposes the automated modeling task into three sub-tasks: first, identifying relationships between entities; second, formalizing the problem into an optimization model using these entities and their relationships; and finally, generating code for testing and execution. Building upon this idea, Li et al. [69] proposed a framework specifically for Mixed-Integer Linear Programming (MILP). In their approach, they employed a fine-tuned model to classify constraints to handle logical constraints and binary variables, which were less explored in previous works. Furthermore, this two-stage method has been successfully applied in various specialized domains. For example, [73] applied it to the field of energy management, requiring the LLM to first extract the five most critical parameters from the problem description before proceeding with automated modeling. Another study [77] applied this approach to travel planning, where the LLM was responsible for extracting user preferences and associated costs, building a complete model, and utilizing a solver to find the optimal travel plan.
Beyond the two-stage framework, researchers have actively explored other prompt-based paradigms to further enhance the performance of large language models in optimization modeling. Two of the most representative paradigms are the multi-agent framework and the interactive framework. The former decomposes the automated modeling task into a series of collaborative expert sub-tasks, which are then completed jointly by multiple LLMs. The latter, on the other hand, aims to alleviate the user’s input burden by not requiring a complete problem description upfront. Instead, the optimization model is progressively refined through human-computer interaction, allowing for incremental additions or removals of constraints, priority adjustments, and other modifications.
In the domain of multi-agent frameworks, the Chain-of-Experts [70] framework proposed by Xiao et al. is a foundational work. They initialized 11 types of experts, including a terminology expert, a modeling expert, a programming expert, and a code review expert, which were coordinated by a conductor agent. This framework dynamically and iteratively selects experts to construct a forward chain-of-thought and incorporates a backward reflection mechanism to handle errors and inconsistencies during the modeling process. Building upon this, subsequent studies have further explored diverse multi-agent collaboration modes [71, 79, 85, 86]. For instance, AhmadiTeshnizi et al. introduced OptiMUS [71] and OptiMUS-0.3 [79], which also use a conductor agent to coordinate modeling, programming, and evaluation tasks. However, this manager reviews the content of each conversational step before assigning the corresponding task to the next agent. However, Wan et al. pointed out that this conductor-orchestrated framework could lead to issues such as an unpredictable workflow and insufficient mathematical accuracy. Therefore, they proposed a cognitively-inspired framework named ORMind [85], which aims to improve the reliability and transparency of solutions while reducing API calls by mimicking the human cognitive process and adopting a structured, predictable workflow alongside a counterfactual reasoning mechanism. Additionally, Mostajabdaveh et al. [72] focused more on the evaluation phase of modeling, replacing the time-consuming process of traditional solver-based verification with peer verification among agents. Similarly, Talebi et al. [84] also centered their research on the evaluation phase, specifically for stochastic optimization problems, by setting up multiple independent reviewer agents to score the modeling results, thereby aiming for a comprehensive assessment of modeling quality.
Both the two-stage and multi-agent frameworks require users to provide a complete problem description before the framework is initiated. However, for many real-world optimization problems, such as conference scheduling and travel planning, user preferences can change dynamically, which necessitates a paradigm that supports continuous interaction between the user and the model. To address this, Zhang et al. proposed the OptLLM [78] framework, which supports both single-shot and interactive inputs. This framework uses an interactive refinement module to help users progressively refine the problem description, while a converter module and a response module handle the subsequent modeling and verification tasks. Lawless et al. [80] further advanced this interactive framework by requiring the LLM to choose from five tasks, such as adding or removing constraints and adjusting priorities, after each user input. Their MeetMeta system not only provides a solution but also explains why certain preferences cannot be met and allows users to make modifications, creating a continuous feedback loop. Subsequent research [81] has focused more on how to translate user priorities into optimization constraints during a dialogue, thereby enabling users to explore trade-offs in tasks like ridesharing coordination.
As research into prompt-based methods has deepened, researchers have begun exploring different avenues to address various existing challenges in LLM-based automated modeling. On the modeling generation side, to tackle the vast and problem-dependent hypothesis space faced by LLMs, researchers have adopted distinct strategies. For instance, Astorga et al. [82] combined LLMs with Monte Carlo Tree Search (MCTS) to efficiently explore this hierarchical and immense hypothesis space, further enhancing search efficiency through symbolic pruning and LLM-based evaluation. In contrast to this search-first strategy, Li et al. [76] employed a generation-first approach. They constrained the LLM’s output using predefined structures, thereby directly avoiding blind searches within the huge hypothesis space. This was complemented by a post-processing verification and correction mechanism to ensure the generated models could be correctly parsed by solvers. On the model verification side, to improve the effectiveness and broaden the scope of validation, Huang et al. [75] proposed an automated evaluation framework based on solver outputs. They also constructed the MAMO benchmark, which extends the NL4OPT benchmark by including systems of ordinary differential equations as well as more complex LP and MILP problems. Conversely, Wang et [105] al. took a different approach, employing a modified Weisfeiler-Lehman (WL) graph isomorphism algorithm to verify model equivalence by transforming optimization models into a bipartite graph structure. They also created the OptiBench benchmark, which contains a large number of LP and MILP problems designed to effectively reflect the performance boundaries of LLMs. Similarly, Zhai et al. [83] introduced the Quasi-Karp equivalence theoretical framework and developed the EquivaMap system, achieving a scalable method for automated equivalence detection through LLM-generated variable mapping functions and lightweight verification.
Despite the significant potential demonstrated by prompt-based frameworks in automated optimization modeling, their technical boundaries remain constrained by the core capabilities of the underlying LLMs. Current mainstream approaches, which rely on closed-source models like the GPT series, face a dual challenge [27]: on one hand, the scarcity of high-quality training data in the optimization modeling domain severely restricts the models’ performance ceiling; on the other hand, using API calls to commercial models may lead to the leakage of sensitive information, posing security risks in high-stakes scenarios such as medical scheduling and defense planning. These fundamental limitations, coupled with the unique formal rigor required for optimization problems, are driving a paradigm shift in research from prompt engineering to learning-based methods. The latter, through fine-tuning open-source models, can achieve more precise mathematical expression generation, more controllable data privacy protection, and more efficient injection of domain knowledge.
IV-B Learning-based Methods
Learning-based methods fundamentally overcome the inherent limitations of prompt engineering by fine-tuning model parameters to internalize optimization modeling knowledge within large language models. In early research, several studies attempted to fine-tune models to enhance specific aspects of automated modeling. For instance, the work [66] combined traditional NLP methods with LLMs to obtain the Ner4OPT model, which is specialized for the NER task. Additionally, Li et al. [69] utilized a fine-tuned model to classify constraints to uncover complex relationships between entities. While these works involved model fine-tuning, their core idea was to serve as auxiliary components within prompt-based two-stage frameworks. Subsequently, the LM4OPT [87] framework proposed by Ahmed et al. significantly improved the optimization modeling capabilities of Llama-2-7b [106] through progressive fine-tuning. However, due to its fine-tuning data being confined to the NL4OPT dataset, its modeling performance still exhibited a substantial gap compared to closed-source models like GPT-4 [107].
Building upon these studies, Huang et al. proposed ORLM [27], a framework featuring a two-stage data synthesis strategy comprising expansion and augmentation. This strategy systematically prepares structured data for model fine-tuning, encompassing three critical components: natural language problem descriptions, formal mathematical models, and executable solver code. Specifically, the expansion stage leverages a small set of seed data and uses GPT-4 to iteratively generate new samples with diverse scenarios and difficulties. The augmentation stage, in turn, further enhances sample variety by employing strategies such as modifying objective functions, altering constraints, and rephrasing problems. During the iterative data synthesis process, the framework also employs regular match correction function to address minor grammatical errors in the code and removes duplicate problems to ensure the integrity and purity of the dataset. Subsequently, they fine-tuned a series of open-source models, such as Mistral-7B [108] and Deepseek-Math-7B-Base [109], using Instruction-Tuning. Comparative tests revealed that the fine-tuned open-source models not only consistently outperformed single-step GPT-4 modeling but also yielded significant performance improvements over classic prompt-based methods like Chain-of-Experts [70] and OptiMUS [71]. The data synthesis-model fine-tuning-evaluation workflow has been widely adopted by subsequent research, with some studies focusing on proposing new data synthesis methods, while others prioritize optimizing the fine-tuning stage.
A subset of learning-based research focuses on the data synthesis component, aiming to address the challenge of data scarcity in optimization modeling. The ReSocratic framework [88] developed by Yang et al. aims to synthesize data in an inverse manner, first synthesizing optimization examples with mathematical formulas and then translating them in reverse into natural language problems, thereby creating a large-scale dataset. Similarly, the OptMATH framework [91] proposed by Lu et al. is a scalable bidirectional data synthesis framework that automatically generates problems with controllable complexity from seed data and uses a forward modeling and rejection sampling mechanism to verify the accuracy of the data pairs. Zhou et al. combined the diversity of forward generation and the reliability of inverse generation from the aforementioned studies to design the DualReflect synthetic data pipeline [95], which is specifically tailored for dynamic programming problems. Beyond focusing on the synthesis methods themselves, other researchers have conducted in-depth studies on the quality of the synthesized content. Wang et al. argued that the data paradigm of methods like ORLM has too large a span from natural language to mathematical models. To address this, they constructed the StructuredOR dataset [90], which aims to solve the problem of missing detailed modeling process annotations (such as variable definitions) in existing datasets, providing more comprehensive annotation information for model learning. Wu et al. focused on generating more complex optimization problems. Their Step-Opt framework [94] systematically increases problem complexity through an iterative problem generation approach and employs a stepwise validation mechanism to ensure data rigor.
Apart from data synthesis, another line of research focuses on fine-tuning methods. Jiang et al. [89] enhanced the model’s accuracy and generalization in problem formalization and solver code generation through multi-instruction fine-tuning, while also introducing model alignment and self-correction mechanisms to mitigate hallucination phenomena. Chen et al. proposed the SIRL framework [93] by incorporating reinforcement learning, which utilizes an external optimization solver as a validator to provide verifiable reward signals for the reinforcement learning process. This approach effectively addresses the hallucination problem in LLM generation, thereby improving the factual correctness and trustworthiness of the generated content. Furthermore, Amarasinghe et al. [92] proposed a method for fine-tuning cost-effective LLMs to tackle specific business challenges such as production scheduling, which demonstrated superior performance in real-world applications.
IV-C Challenges
With the growing application of LLMs in optimization, increasing efforts have been made to bridge natural language descriptions of optimization problems with formal mathematical models. Frameworks such as Chain-of-Experts [70] and ORLM [27] demonstrate that LLMs can not only complete modeling tasks efficiently but also uncover implicit constraints that may be overlooked by human experts, thus introducing a new paradigm for optimization modeling. Nevertheless, existing research still faces several key challenges:
  • Prompt-based methods are widely adopted due to their ease of deployment and the absence of large annotated datasets. However, their reliance on closed-source LLM APIs raises serious concerns regarding data security and privacy, especially in industrial settings involving sensitive information. Moreover, even advanced models such as GPT-4 exhibit limitations in mathematical reasoning, leading to issues of accuracy and robustness on fundamental modeling tasks [27]. Hence, enhancing the intrinsic reasoning capability of LLMs remains a primary challenge for this paradigm.
  • Learning-based methods such as ReSocratic [88] and OptMATH [91] have investigated synthetic data generation to construct large-scale optimization datasets. Yet, studies suggest that comparable performance can sometimes be achieved using only a fraction of the dataset, pointing to diminishing returns from data scaling [110]. Consequently, an open problem is how to efficiently generate high-quality, representative datasets that maximize learning efficiency. Furthermore, as this line of research is still at an early stage, the number of existing studies remains limited, and broader exploration is needed to fully exploit the potential of learning-based modeling.
V LLMs for Optimization Solving
Optimization solving is a core stage of the optimization workflow, and decades of research have produced a wide range of high-quality algorithms. In recent years, machine learning methods have been increasingly integrated into optimization to improve performance, and the advent of LLMs has further propelled this research into a new phase. In this section, we categorize existing work into three paradigms: LLMs as optimizers, low-level LLM-assisted optimization algorithms, and high-level LLM-assisted optimization algorithms. Section V-A introduces LLMs as optimizers, which directly employ LLMs to solve problems without relying on traditional frameworks (see Fig. 4). Sections V-B and V-C discuss how LLMs can enhance optimization algorithms at the component level and at the orchestration/design level, respectively (illustrated in Figs. 5 and 6). Section V-D then outlines the key challenges in this area. Representative studies are summarized in Table III.
TABLE III: Representative works on LLMs for optimization solving, categorized into LLMs as optimizers, low-level LLMs for optimization algorithms and high-level LLMs for optimization algorithms.
Method Venue Type Technical Summary LLMs as Optimizers OPRO [28] ICLR, 2023 Prompt-based Iteratively refine through optimization trajectories and problem descriptions. toLLM [111] KDD, 2023 Prompt-based Design four canonical tasks to evaluate the performance boundaries of LLMs. EvoLLM [112] GECCO, 2024 Prompt-based Replace traditional optimization trajectories with candidate solution quality ranking. MLLMO [113] MCII, 2024 Prompt-based Utilize MLLMs to jointly process problem descriptions and map visualizations for CVRPs. POM [114] NeurIPS, 2024 Learning-based Pre-train a general-purpose, zero-shot black-box optimization foundation model. OPTO [115] NeurIPS, 2024 Prompt-based Replace traditional optimization trajectories with rich, structured execution traces. VRMA [116] arXiv, 2024 Prompt-based Utilize MLLMs to process 2D planar point distribution maps as input. ROPRO [117] ACL, 2024 Prompt-based Discuss the model dependency of OPRO and identify its limitations on small-scale models. BBOLLM [29] arXiv, 2024 Prompt-based Conduct evaluation of LLMs on both discrete and continuous black-box optimization problems. LLMS [118] arXiv, 2024 Learning-based Fine-tune LLMs with instruction-solution pairs for scheduling problems. ECLIPSE [119] NAACL, 2024 Prompt-based Apply iterative optimization to jailbreaking attack strategies. LLOME [120] arXiv, 2024 Learning-based Integrate preference learning to train LLMs for satisfying complex biophysical constraints. LLMDSM [121] arXiv, 2024 Prompt-based Apply iterative optimization to design structure matrix sequencing. LMCO [122] WCL, 2024 Prompt-based Apply iterative optimization to wireless network design. MGSCO [123] arXiv, 2025 Prompt-based Process visual representations of abstract graph structures using MLLMs. ORFS [124] arXiv, 2025 Prompt-based Apply iterative optimization to automated parameter tuning in chip design. Low-level LLMs for Optmization Algorithms LMX [125] TELO, 2023 Operators Leverage LLMs as intelligent operators for textual genome crossover and recombination. GPT-NAS [126] arXiv, 2023 Initialization Utilize LLMs for NAS initialization with prior knowledge. LMEA [127] CEC, 2023 Operators Employ LLMs as crossover, mutation, and selection operators to guide EAs. LLM-PP [128] arXiv, 2023 Initialization Utilize LLMs as performance predictors to assist initialization processes. LMOEA [129] EMO, 2023 Operators Empower MOEA/D with LLMs through zero-shot prompting as search operators in multi-objective optimization. OFPLLM [130] AI, 2023 Initialization Assist non-expert users in initializing financial plans. LEO [131] Neucom, 2024 Operators Utilize LLMs to guide individuals from dual pools for exploration and exploitation. LLM-GA [132] BiB, 2024 Initialization Initialize genetic algorithms with LLMs to generate high-quality mutant pools for enzyme design. GE [133] GECCO, 2024 Operators Enhance the creativity and diversity of the LLM by introducing different role-based prompts for NAS. LLMAES [134] ICIC, 2024 Operators Generate minimal solutions within populations using LLMs to reduce interaction costs. LLMTES [135] arXiv, 2024 Configuration Apply LLM-based feedback loops to sequentially optimize evolution strategies through static tuning. LAEA [136] SWEVO, 2024 Evaluation Transform model-assisted selection tasks into classification and regression problems. LICO [137] arXiv, 2024 Evaluation Apply LLMs as surrogate models in molecular science applications. LLMSM [138] CAI, 2024 Evaluation Develop a collaborative framework using LLMs for model selection and training in engineering optimization. LLMCES [139] GECCO, 2024 Configuration Implement step-size control for evolution strategies using OPRO-like mechanisms. AKFLLM [140] arXiv, 2024 Operators Utilize LLMs to assist mutation and other generative stages in evolutionary multi-task optimization. LLMAMO [141] arXiv, 2024 Operators Invoke LLMs for elite solution generation only upon insufficient population improvement. LTC [142] ESANN, 2024 Configuration Apply LLM-based feedback loops to sequentially optimize CMA-ES through dynamic control. LLMEVO [143] arXiv, 2025 Operators Validate LLM effectiveness in selection, crossover, and mutation phases while noting limitations in initialization. LLM-GE [144] GECCO, 2025 Operators Utilize LLMs as crossover and mutation operators to optimize YOLO architectures for object detection. LAOS [145] GECCO, 2025 Configuration Utilize state features to replace optimization trajectories for guiding LLM-based operator selection. PAIR [146] arXiv, 2025 Operators Focus on utilizing LLMs as selection operators to enhance LMEA. LLMMS [147] arXiv, 2025 Evaluation Utilize LLMs as meta-Surrogates to facilitate cross-task knowledge sharing through token sequence representations. LMPSO [148] arXiv, 2025 Operators Simulate PSO evolutionary processes using LLMs and refine LMEA for specific algorithms. LLM-SAEA [149] arXiv, 2025 Evaluation Dynamically select appropriate proxy models and infill sampling criteria using LLMs. High-level LLMs for Optmization Algorithms LLMGM [150] GECCO, 2023 Single-step Decompose and recombine six existing metaheuristic algorithms to generate novel hybrid algorithms in single rounds. FunSearch [151] Nature, 2023 Iterative Generate novel code snippets via LLMs and guide EAs to search through function space. AEL [152] arXiv, 2023 Iterative Enhance FunSearch by integrating heuristic principles, demonstrating significant improvements on TSP instances. AS-LLM [153] IJCAI, 2023 Selection Automatically extract high-dimensional algorithm features from code or descriptive text to facilitate algorithm selection. EoH [154] ICML, 2024 Iterative Implement heuristic-code coevolution via preset prompt pairs for algorithm crossover and mutation. ReEvo [155] NeurIPS, 2024 Iterative Utilize LLMs as hyper-heuristic algorithms with reflective evolution mechanisms to guide search processes. AutoSAT [156] arXiv, 2024 Iterative Employ multiple heuristic strategies to guide LLMs in algorithm generation. LLaMEA [157] TEVC, 2024 Iterative Utilize LLMs to guide EAs in mutation, selection, and other operations for generating advanced heuristics. L-AutoDA [158] GECCO, 2024 Iterative Apply AEL to generate adversarial attack algorithms in cybersecurity. MOELLM [159] TEVC, 2024 Iterative Introduce robust testing modules and dynamic selection strategies to generate multi-objective optimization algorithms. LLMEPS [160] PPSN, 2024 Iterative Develop a baseline algorithm for automated algorithm design using EoH and ReEvo methodologies. EvolCAF [161] PPSN, 2024 Iterative Apply EoH to evolve and generate cost-aware acquisition functions for Bayesian optimization. TS-EoH [162] arXiv, 2024 Iterative Apply EoH to generate algorithms for edge server task scheduling. MEoH [163] AAAI, 2024 Iterative Extend EoH to multi-objective versions, incorporating other metrics for comprehensive algorithm evaluation. LLaMEA-HPO [164] TELO, 2024 Iterative Integrate hyperparameter optimization into LLaMEA’s iterative cycle to enhance overall efficiency. CMLLM [165] EvoApps, 2024 Iterative Develop dynamic prompt mechanisms to control LLM mutation for enhancing LLaMEA performance. HSEvo [166] AAAI, 2024 Iterative Combine harmony search with GAs and optimize algorithm portfolio quality/diversity through diversity metrics. LLM4AD [167] arXiv, 2024 Iterative Develop an EoH-based platform to accelerate heuristic algorithm design. MCTS-AHD [168] arXiv, 2025 Iterative Employ MCTS to explore heuristic spaces, organizing generated heuristics into tree structures. CEG [169] GECCO, 2025 Iterative Develop methodologies to analyze LLM-generated code and its evolutionary patterns during optimization processes. OPSLAD [170] GECCO, 2025 Iterative Apply LLaMEA to photonic structure optimization in industrial design. BLADE [171] GECCO, 2025 Iterative Establish a standardized and reproducible evaluation framework for LLM-driven algorithm generation methods. AutoHD [172] arXiv, 2025 Iterative Propose an LLM-guided heuristic discovery strategy for complex planning tasks. LAMA [173] TEVC, 2025 Iterative Propose an automatic memetic algorithm enhanced by a heuristic designed with the assistance of LLMs. LAS [174] arXiv, 2025 Iterative Analyze fitness landscapes of LLM-assisted algorithm search processes. LLaMEA-BO [175] arXiv, 2025 Iterative Apply LLaMEA to automatically generate Bayesian optimization algorithms. InstSpecHH [176] arXiv, 2025 Selection Leverage LLMs’ semantic comprehension and reasoning capabilities for context-aware algorithm selection. BSALMD [177] arXiv, 2025 Iterative Conduct behavioral space analysis to examine algorithmic evolution trajectories and behaviors. FLAAD [178] arXiv, 2025 Iterative Fine-tune LLMs with diversity-aware ranked sampling for algorithm generation. EoH-S [179] arXiv, 2025 Iterative Extend EoH to evolve and generate complementary algorithm portfolios.
Figure 4: Illustration of LLMs as optimizers. This paradigm primarily relies on interactive prompting for optimization, with a few studies incorporating fine-tuning and pre-training for enhanced performance.
V-A LLMs as Optimizers
LLMs possess strong in-context learning capabilities, enabling them to solve optimization problems through iterative prompting. A pioneering work in this direction is the OPRO framework proposed by Yang et al. [28], which formulates optimization tasks in natural language and iteratively refines solutions based on historical trajectories. By progressively improving solution quality, OPRO was the first to demonstrate the feasibility of using LLMs as general-purpose optimizers, inspiring subsequent research across diverse domains. Building on this paradigm, Guo et al. [111] benchmarked LLMs on four classical optimization tasks i.e., gradient descent, hill climbing, grid search, and black-box optimization, where the model iteratively generated and evaluated solutions through interactive prompting. To ensure fair and consistent evaluation, they also introduced three task-agnostic metrics designed to reduce sensitivity to test sample variations. Beyond benchmarking, OPRO has been extended to multiple application domains. Jiang et al. [121] applied it to combinatorial optimization for structural matrix ordering; Qiu et al. [122] developed LMCO for wireless network design; Ghose et al. [124] employed it for parameter tuning in open-source hardware workflows; and Jiang et al. [119] adapted the paradigm to jailbreak attack analysis, treating optimizable suffixes and harmfulness scores as part of the optimization trajectory.
In contrast to OPRO’s direct use of solution–feedback trajectories, subsequent research has explored alternative forms of iteration to guide LLMs in optimization tasks. Lange et al. [112] sorted discretized candidate solutions from best to worst and provided the ordered sequence as context to the LLM. This reframes the task from pure solution generation to sequence recognition and reasoning, leveraging the model’s ability to capture ordering patterns. Cheng et al. [115] proposed the OPTO framework, which feeds LLMs with structured execution traces (e.g., intermediate variable values, function calls, and runtime outputs) to simulate backpropagation-like optimization and improve task understanding. Beyond textual iteration, multimodal extensions have further broadened the paradigm by incorporating visual information. Huang et al. [113] addressed vehicle routing problems using both textual descriptions and map images to mimic human spatial cognition. Elhenawy et al. [116] solved the classic traveling salesman problem by combining scatter plot inputs with a multi-agent collaboration mechanism for path search. Zhao et al. [123] extended this idea to graph-structured problems by converting abstract graphs into visual images, enabling multimodal LLMs to capture topological features in a manner more akin to human reasoning. Collectively, these studies demonstrate that extending OPRO with sequence-based, trace-based, or multimodal inputs can substantially enrich the optimization capabilities of LLMs.
Despite the progress of using LLMs directly as optimizers, several inherent drawbacks have been identified. Zhang et al. [117] highlighted the strong model dependency of OPRO, noting that smaller models, with limited reasoning capacity, struggle to exploit iterative histories and generate high-quality solutions. Huang et al. [29] systematically evaluated LLMs on both discrete and continuous black-box optimization problems, confirming their strengths in symbolic and text-based discrete tasks but also exposing their weaknesses in purely numerical settings, such as continuous optimization.
Figure 5: Illustration of low-level LLMs for optimization algorithms. LLMs can be applied at various stages within EAs, including initialization, evolutionary operators, algorithm configuration, and fitness evaluation.
Based on these limitations, researchers have begun to explore alternative avenues for applying LLMs in optimization. One line of work attributes the drawbacks to prompt engineering and the inherent constraints of current models, and thus seeks to improve optimization capability through fine-tuning or training. Abgaryan et al. [118] fine-tuned an LLM to output solutions in a single step, bypassing multi-round iteration by training on instruction–solution pairs. Li et al. [114] introduced POM, a pretrained optimizer model that generalizes across black-box problems in a zero-shot manner. Chen et al. [120] proposed LLOME, a two-level optimization framework that incorporates preference learning to generate biologically valid sequences under complex biophysical constraints. Meanwhile, an increasing body of research suggests that LLMs may be more effective when used to assist rather than replace traditional optimization algorithms. This perspective motivates the next paradigm of low-level LLM-assisted optimization, where LLMs are embedded into existing frameworks to enhance their performance.
V-B Low-level LLMs for Optimization Algorithms
There is a natural complementarity between the collective intelligence of EAs and the knowledge-driven reasoning of LLMs [32]. While EAs excel at heuristic exploration through population-based search, LLMs contribute powerful semantic understanding and generative capabilities that can guide this process, thereby improving both efficiency and solution quality. This collaborative paradigm, which embeds LLMs as components within EAs, defines the direction of low-level assistance in optimization. In the following, we systematically review research in this area, focusing on four key stages of EAs where LLMs have been applied: initialization, evolutionary operators, algorithm configuration, and evaluation.
V-B1 Initialization
In the optimization process, the quality of the initial solution strongly influences both convergence efficiency and final performance. By leveraging their rich background knowledge, LLMs can provide high-quality initial solutions that embed prior information, thereby narrowing the search space and establishing a stronger foundation for subsequent iterations. Representative applications have demonstrated this potential across diverse domains. In Neural Architecture Search (NAS), Yu et al. [126] employed LLMs to propose plausible architectural components, introducing valuable prior knowledge before the search began. Jawahar et al. [128] used LLMs as predictors to guide the initialization phase of NAS more efficiently. Beyond NAS, Teukam et al. [132] applied LLM-based initialization in genetic algorithms for bioengineering, generating high-quality mutant libraries to accelerate enzyme design. In financial planning, De et al. [130] demonstrated that LLMs can generate viable initial portfolios for optimization tasks. Despite these advances, challenges remain. Zhao et al. [143] highlighted the high computational cost of using LLMs for initialization, as well as their difficulty in consistently satisfying strict constraints and generating feasible solutions, particularly in large-scale or highly complex problem settings.
V-B2 Evolutionary Operators
Evolutionary operators are the core drivers of EAs. Before the advent of LLMs, researchers explored the use of neural networks, such as attention mechanisms and feed-forward networks [180], to simulate the behavior of individual operators or even entire algorithms. With the rise of iterative prompting approaches such as OPRO [28], attention has shifted toward employing LLMs directly to simulate operator behavior. This paradigm enables more adaptive search during optimization, as LLMs can exploit problem and solution characteristics without additional training, thereby surpassing the performance of conventional operators in some settings.
Several frameworks exemplify this direction. The LMX framework proposed by Meyerson et al. [125] employs few-shot prompting to use LLMs as intelligent operators for crossover and recombination of text-based genomes, generating semantically coherent variants without training. Liu et al. [127] extended this idea through the LMEA framework, where LLMs perform selection, crossover, and mutation under prompt guidance, coupled with a temperature self-adaptive mechanism to balance exploration and exploitation. Other studies have targeted specific operators or algorithms: PAIR [146] focuses on selection, LMPSO [148] adapts LLMs to PSO, and LEO [131] explores LLM-driven strategies for balancing exploration and exploitation.
This line of work has also expanded into multi-objective optimization. Liu et al. [129] were the first to empower MOEA/D [181] with zero-shot prompting, allowing LLMs to act as search operators and inspiring the design of more efficient white-box operators. To reduce interaction costs, Wang et al. [134] proposed a hybrid scheme in which LLMs generate only 10% of the population, while the rest are produced by traditional operators. Similarly, Liu et al. [141] called upon LLMs only when population improvement was insufficient, while relying on NSGA-II [182] otherwise. These hybrid approaches highlight the potential of selectively integrating LLMs to balance performance gains with computational efficiency.
V-B3 Algorithm Configuration
The performance of EAs is strongly influenced by hyperparameter and operator configurations. Traditionally, configuration has been studied through static tuning [183] and dynamic control [184]. In recent years, dynamic configuration has become particularly prominent, largely driven by RL methods [185, 186, 187], given the natural alignment between the Markov Decision Process (MDP) framework [188] and adaptive configuration. However, RL-based approaches usually require problem-specific training, leading to high computational costs and limited generalization. Leveraging their reasoning abilities, LLMs provide a promising alternative by analyzing optimization features and recommending adaptive configurations without extensive retraining.
Several representative works illustrate this direction. Kramer et al. introduced an LLM-based feedback loop for both static tuning [135] and dynamic control [142] of evolutionary strategies. Custode et al. [139] applied an OPRO-based mechanism, feeding optimization trajectories back to the LLM for step-size control, and additionally prompted the model to justify its parameter updates. Zhang et al. [145] addressed the redundancy of trajectory-based feedback by proposing a meta-prompt built on features from the solution space, decision space, and optimization process. They further integrated traditional adaptive control and prior optimization experience to enable LLM-guided adaptive operator selection.
It is worth noting that some algorithm generation frameworks also involve hyperparameter tuning. For example, EoH [154] employs prompt templates for parameter tuning during heuristic mutation, while LLaMEA [157] integrates hyperparameter optimization to reduce iteration costs. However, the primary contribution of these works lies in algorithm generation rather than EA configuration, and thus they fall outside the strict scope of configuration research.
V-B4 Evaluation
In computationally expensive optimization problems, surrogate models are widely used to approximate costly real-world evaluations and alleviate computational bottlenecks. Recently, LLMs have been explored as powerful tools for surrogate modeling, leveraging their pattern recognition and reasoning capabilities to improve evaluation efficiency.
A pioneering study by Hao et al. [136] introduced an LLM-based surrogate that reframes model-assisted selection as classification and regression, enabling evaluation of new solutions from historical data without additional training. Building on this idea, Nguyen et al. [137] developed the LICO framework to address data scarcity in domains such as molecular optimization, where direct LLM application is challenging. Beyond acting as surrogates, LLMs have also been employed for managing surrogate-assisted algorithms. Rios et al. [138] incorporated LLMs into the selection and training of surrogate models for engineering optimization. Xie et al. [149] further proposed LLM-SAEA, a framework that uses collaborative experts to dynamically select surrogate models and infill criteria, reducing reliance on extensive domain expertise.
More recently, Zhang et al. [147] extended this line of research to multi-task optimization, treating LLMs as meta-surrogates capable of facilitating knowledge transfer across tasks through unified token sequence representations. These studies collectively highlight the potential of LLMs not only as direct surrogates but also as adaptive coordinators in surrogate-assisted optimization, though further work is required to assess their robustness and scalability in large-scale applications.
V-C High-level LLMs for Optimization Algorithms
In contrast to low-level assistance, which targets micro-level interventions within EAs, high-level assistance operates at the macro level. Its goal is to enhance solving performance by leveraging LLMs to intelligently select among existing algorithms or to generate entirely new ones, thereby exploiting performance synergies across methods. In the following, we systematically examine two major forms of high-level assistance: algorithm selection and algorithm generation.
V-C1 Algorithm Selection
The goal of algorithm selection is to identify the most suitable algorithm from a portfolio for each problem instance. This problem arises from the significant performance disparities among algorithms, which often exhibit different optimization behaviors and application domains [7]. Traditionally, algorithm selection is formulated as a machine learning task, where a selector is trained on problem instances and their features to predict the best algorithm for unseen cases [189]. The process typically involves two stages: feature extraction and selector construction. The quality of extracted features directly affects the quality of training samples and thus the performance of the selector. Commonly used features include basic statistical descriptors of the problem [190] as well as optimization characteristics obtained from algorithm runs [191]. Selector construction can then be approached through classification [192], regression [193], or hybrid models [194], offering a wide range of technical solutions for this task.
Figure 6: Illustration of high-level LLM-assisted optimization algorithms. Algorithm selection involves two key stages: feature extraction and selector construction. Algorithm generation has evolved from single-step to iterative generation, reflecting the shift toward more adaptive and context-aware design.
Based on this workflow, recent studies have applied LLMs to both feature extraction and selector construction. Wu et al. introduced the AS-LLM framework [153], the first to bring LLMs into algorithm selection. AS-LLM leverages the code understanding capabilities of LLMs to automatically extract high-dimensional algorithm features from code or descriptive text. Its core representation module combines LLM-extracted features with a feature selection mechanism to identify the most relevant subset. The resulting algorithm features and problem features are then processed in separate networks to avoid premature interaction, and a similarity calculation module determines the final choice by evaluating the match between problems and algorithms. In contrast, Zhang et al. proposed InstSpecHH [176], which emphasizes selector construction. After a Euclidean distance–based pre-selection step narrows the candidate algorithms, the natural language description of the problem instance and the pre-selected algorithms are fed into the LLM. This approach exploits the LLM’s semantic reasoning ability to achieve more context-aware selection. It is worth noting that some related works involve solver selection [71, 79], where LLMs are integrated into broader workflows such as optimization modeling or end-to-end pipeline design. While relevant, these works differ from the classical definition of algorithm selection and are generally not categorized as such.
V-C2 Algorithm Generation
In contrast to algorithm configuration and selection, algorithm generation aims to automatically design optimization algorithms from problem characteristics without relying on pre-defined component or algorithm pools. Reinforcement learning–based approaches have made progress in this direction, but they still depend on manually specified component or symbol sets [35]. For instance, the GSF framework [195] can generate algorithms within a generic EA template, yet each component pool must be pre-specified. Chen et al. proposed SYMBOL [17], which removed the component pool constraint but continued to rely on a manually defined symbol set. LLMs, with their broad knowledge bases and strong generative capabilities, offer a promising way to overcome these limitations. By directly producing algorithmic structures and logic, LLMs bypass the need for pre-specified pools, making algorithm generation one of the most active branches of LLM-assisted optimization. Research in this area has evolved from early single-step generation toward more advanced iterative generation frameworks [153].
Single-step generation represents the initial exploration of this field. Pluhacek et al. [150] used GPT-4 to decompose and recombine six metaheuristic algorithms, creating new hybrid variants. Zhong et al. [196] combined GPT-3.5 with a prompt engineering framework to generate ZSO, a novel metaheuristic, in a single step. These studies demonstrated the feasibility of LLM-driven algorithm generation but also revealed the limitations of single-step methods, which depend heavily on the model’s inherent capabilities. Consequently, current research has shifted toward iterative generation approaches, which provide greater adaptability and robustness.
Iterative generation has since become the dominant paradigm, with research progressing rapidly from foundational studies to real-world applications. A pioneering work in this direction is Funsearch [151], which integrates LLMs with evolutionary algorithms to iteratively generate program snippets in function space. Building on this idea, Liu et al. proposed AEL [152] and EoH [154], where heuristic concepts co-evolve with code, achieving superior performance on combinatorial optimization benchmarks. To support this line of work, they developed the LLM4AD platform [167] and analyzed the multimodal and rugged fitness landscape of LLM-assisted algorithm search [174]. More recently, they fine-tuned LLMs specifically for algorithm design, introducing diversity-aware ranking, sampling strategies, and direct preference optimization, demonstrating that fine-tuned models can significantly outperform generic ones [178]. The EoH framework has since been extended into MEoH [163] for multi-objective heuristic search and EoH-S [179] for co-evolving complementary sets of algorithms, with applications in edge server task scheduling [162], Bayesian optimization [161], and adversarial attack design [158].
In parallel, Van Stein and Bäck developed the LLaMEA series [157, 197], which integrates LLMs into evolutionary algorithms for tasks such as mutation and selection, enabling the generation of advanced heuristics. To better characterize the generation process, they introduced code evolution graphs [169] and behavior space analysis [177], and built the BLADE framework [171] to provide standardized evaluation for LLM-driven algorithm generation. Extensions such as LLaMEA-HPO [164] incorporated hyperparameter optimization to improve efficiency, while applications have spanned Bayesian optimization [175] and photonic structure design [170].
Beyond these two main lines, several studies have explored novel search paradigms. Ye et al. proposed ReEvo [155], which treats the LLM as a hyper-heuristic and introduces a reflective evolution mechanism. Zheng et al. developed MCTS-AHD [168], employing Monte Carlo Tree Search to explore heuristic space more thoroughly and avoid local optima. Pham et al. proposed HSEvo [166], combining harmony search with genetic algorithms and introducing diversity metrics to balance exploration and exploitation. Together, these efforts highlight the shift from single-step demonstrations toward robust iterative frameworks that can adaptively generate increasingly sophisticated optimization algorithms.
V-D Challenges
When applying LLMs to optimization solving, methods at different levels exhibit distinct characteristics. The direct use of LLMs as optimizers, exemplified by OPRO [28], marked the initial integration of LLMs with optimization. Low-level approaches, such as LMEA [127] and LMX [125], have been deployed in domains like NAS and molecular discovery. High-level methods for algorithm generation, including EoH [154] and LLaMEA [157], currently represent the most active research frontier. While these studies have advanced the synergy between LLMs and optimization, several bottlenecks remain:
  • LLMs as Optimizers: Direct, trajectory-based optimization with LLMs often struggles to outperform classical algorithms [29] and performs poorly with smaller models that lack sufficient reasoning ability [117]. Moreover, trajectory-based iteration is difficult to extend to population-based optimization, since the explosion of trajectory data easily exceeds the LLM’s context capacity. Future work should explore alternative iterative forms that replace raw trajectories with more structured representations of optimization history.
  • Low-level LLMs for Optimization Algorithms: While the use of LLMs as evolutionary operators has shown promise in engineering, biology, and chemistry, research on initialization and algorithm configuration remains nascent. LLMs are difficult to apply to data-intensive tasks like large-scale initialization [143], and they often display conservative behavior in parameter control [139], failing to provide sufficient exploration. The underlying challenge is the vastness of the decision space. Future directions should focus on reducing decision-making pressure while preserving accuracy.
  • High-level LLMs for Optimization Algorithms: Major issues arise in both computational cost and innovation. Frameworks such as EoH [154] and LLaMEA [157] treat algorithms as individuals in an EA, which requires frequent LLM calls and costly evaluations during evolution. In terms of innovation, current methods mainly recombine existing strategies and remain limited in generating fundamentally novel approaches. Achieving scalable, low-cost search and fostering true algorithmic creativity are key open challenges for this line of research.
VI Applications
Beyond classical numerical and combinatorial optimization problems, LLMs have been applied to optimization tasks across diverse interdisciplinary domains. This section highlights representative applications in computer science, natural sciences, and engineering. Section VI-A discusses applications in machine learning and computer security. Section VI-B covers applications in chemistry, biology, and physics. Section VI-C focuses on modern engineering and industrial use cases.
VI-A Computer Science
In computer science, LLM-assisted optimization has been explored in multiple areas, with neural architecture search [198] and computer security [158] standing out as representative examples. Other emerging applications [199, 200, 201, 202] further illustrate the breadth of this direction.
NAS is a core problem in machine learning, where EAs have been widely applied [203]. Integrating LLMs offers new opportunities. While a few studies have attempted to treat LLMs as direct optimizers [204], most research emphasizes low-level assistance. Chen et al. [198] pioneered the use of LLMs as crossover and mutation operators within EAs, where crossover generates new architectures through few-shot prompting and mutation introduces diversity via mixed-temperature sampling. Building on this, Nasir [205] applied the approach to Quality-Diversity algorithms, achieving more efficient exploration with a dual-archive mechanism. Morris et al. [133] enhanced search diversity using role-based prompts, while Mo et al. [206] extended the paradigm to graph neural networks. In addition, LLMs have been explored for initialization: Yu et al. [126] leveraged pre-training, and Jawahar et al. [128] employed few-shot learning to provide high-quality initial suggestions.
In security-related optimization, LLMs have been used both as standalone optimizers and as auxiliary operators. Jiang et al. [119] applied LLMs directly to iteratively generate jailbreak attack suffixes with self-reflection. Complementary work explored low-level assistance [207, 208], where LLMs act as heuristic operators to guide jailbreak prompt optimization, improving both efficiency and transferability. Guo et al. [158] investigated LLM-enabled optimization of the generation function responsible for producing adversarial samples, supporting continuous innovation and evaluation of attack strategies.
Beyond NAS and security, LLM-based optimization has also been applied to emerging tasks. Wang et al. [199] treated data augmentation as an optimization problem, representing augmentation strategies as editable natural language instructions. By simulating crossover and mutation, LLMs intelligently explored strategies tailored to long-tailed distributions. Li et al. [200] focused on GPU computing, where LLMs iteratively generated and optimized CUDA kernel code, achieving systematic performance improvement without human intervention.
VI-B Natural Sciences
In the natural sciences, including chemistry [209], biology [210], and physics [211], many fundamental problems can be cast as complex optimization tasks. These problems are often characterized by high-dimensional search spaces, implicit domain knowledge, and computationally expensive simulations. The reasoning and generative capabilities of LLMs provide a novel approach to addressing these challenges.
In chemistry, LLM-assisted optimization has been widely explored in molecular discovery [209]. Most studies emphasize low-level methods that embed LLMs into EAs. For example, Wang et al. [212] integrated LLMs enhanced with chemical knowledge to perform crossover and mutation, guiding molecular design more effectively. Guevorguian et al. [213] replaced genetic operators with fine-tuned LLMs and introduced a prophetic model to avoid stagnation, later extending this idea to multi-objective optimization for molecular discovery [214]. Beyond operator-level assistance, some works explored surrogate modeling, such as Nguyen et al. [137], who applied LLMs to mitigate data scarcity in molecular optimization.
Similar to chemistry, biological applications often adopt low-level paradigms. The LLM-GA framework [210] uses LLMs for population initialization, generating high-quality mutant libraries for enzyme design. Tran et al. [215] and Wang et al. [216] employed LLMs as intelligent mutation or crossover operators for protein design. Beyond operator-level integration, Chen et al. [217, 120] developed a two-level optimizer that iteratively modifies biological sequences, while Lv et al. [218] proposed a workflow spanning modeling and solving, translating natural language descriptions into protein function predictions and optimizing sequences iteratively.
Applications in physics are more diverse, spanning fluid dynamics [211, 219], semiconductor physics [220], and statistical physics [221, 222]. At the modeling level, Du et al. [221] used GPT-3.5 to derive partial differential equations from data, while Zhang et al. [211] applied LLMs to turbulence closure modeling. Ma et al. [223] proposed a holistic workflow that integrates both modeling and solving. At the solving level, Li et al. [220] incorporated LLMs into optimization workflows to generate and debug code for optimizing laser parameters, while Zhang et al. [219] directly used LLMs as optimizers to generate parameter combinations for fluid dynamics.
VI-C Engineering and Industry
In engineering and industry, LLMs are increasingly applied to complex optimization problems. Representative domains include wireless networks and communications [122], industrial design [170], and edge computing and scheduling [162], with wireless networks and industrial design being the most active areas.
In wireless networks, LLMs have been employed across the full workflow from modeling to solving. On the modeling side, Peng et al. proposed LLM-OptiRA [224], which automatically detects non-convex components in wireless resource allocation problems and reformulates them into solvable forms. Wen et al. [225] further introduced Retrieval-Augmented Generation (RAG) to incorporate external expert knowledge, improving modeling accuracy. On the solving side, Qiu et al. [122] directly used LLMs as optimizers for combinatorial tasks such as access point placement, with subsequent work extending this paradigm to resource allocation [226], power control [227], and multi-UAV deployment [228]. Li et al. [229] explored low-level assistance by embedding LLMs into a multi-objective optimization framework for ISAC systems in UAV networks. More recently, Hou et al. proposed the LHS framework [230], combining heuristic recommendation, enhanced initialization, and iterative refinement.
In industrial design, LLMs have been used for both algorithm configuration and generation. Ghose et al. [124] designed an LLM-based agent for chip design, enabling parameter tuning to optimize performance, power, and area. Jiang et al. [121] applied LLMs as optimizers for design structure matrices, capturing interdependencies within engineering systems. Yin et al. [170] explored algorithm generation in optical design, applying the LLaMEA framework [157] to discover novel algorithms for photonic structure design. These examples highlight how LLMs introduce fundamentally new approaches to long-standing industrial optimization problems.
Beyond networks and design, LLMs are also applied in edge computing and scheduling. For instance, Yatong et al. [162] leveraged LLMs for task scheduling in edge server environments, demonstrating their utility in distributed system optimization.
VII Vision for the Field
The rapid emergence of LLMs for optimization has produced diverse methods across modeling, solving, and application domains. Yet, despite remarkable progress, the field is still in its infancy. Looking forward, we highlight three paradigm shifts that will shape its next stage of development: bridging modeling and solving, transitioning from static to dynamic methods, and building a broader agentic optimization ecosystem.
VII-A Bridging Modeling and Solving
Although LLMs have been successfully applied to both optimization modeling and solving, these domains remain largely disjoint. Frameworks such as OptiMUS [71, 79] and ORLM [27] begin to connect the pipeline: OptiMUS employs an agent-based conductor for modeling and programming, while ORLM fine-tunes LLMs to generate mathematical models and solver code. Yet in all cases, the crucial solving step still depends on external, traditional solvers. This confines LLMs to the front-end, without incorporating advances in LLM-assisted solving such as adaptive configuration, selection, or algorithm generation.
A critical future direction is to close this gap by constructing end-to-end LLM-driven workflows, where problem understanding, model formulation, algorithm design, execution, and evaluation form a self-sufficient loop. Such integration would move beyond reliance on external solvers, harnessing the generative and reasoning power of LLMs to create holistic, closed-loop optimization frameworks capable of autonomously tackling complex problems.
VII-B Transitioning from Static to Dynamic Methods
Current research in LLMs for optimization spans tasks such as optimization modeling, algorithm configuration, selection, and generation. Yet most approaches remain static, relying on fixed prompts, predefined operators, or single-shot algorithm designs. By contrast, the reinforcement learning community has developed rich paradigms for dynamic configuration [47, 231], selection [232], and generation [17], where algorithm behavior adapts continuously to real-time optimization features. For example, dynamic selection methods can exploit the complementary strengths of multiple algorithms through adaptive scheduling, while dynamic generation can synthesize novel operators tailored to evolving search states.
In LLM-based optimization, only a handful of works explore dynamic behavior, such as interactive modeling [80] or adaptive parameter control [139, 145]. Critically, dynamic algorithm selection and generation remain almost untouched. We envision a decisive shift toward methods that not only adapt in real time but also enable the self-evolution of algorithms: frameworks in which LLMs can continually refine, recombine, and reinvent optimization strategies based on feedback from the search process itself. Such self-evolving systems would transform LLMs from passive assistants into adaptive entities capable of lifelong learning, bridging optimization with open-ended algorithmic innovation.
VII-C Toward an Agentic Ecosystem for Optimization
The role of LLMs in optimization should ultimately be conceived not as isolated engines, but as autonomous agents within a broader agentic ecosystem for optimization. In such systems, LLMs function as specialized agents collaborating with others (e.g., human experts, domain-specific solvers, and simulators) through communication, negotiation, and coordination. Early studies already suggest this potential: LLMs can capture tacit expertise, such as driver knowledge in routing [233], domain constraints in scheduling [234], and user preferences in traveling [235], making it actionable within optimization workflows [70, 236, 80, 202]. Similarly, agentic interfaces have enhanced accessibility and trust, where LLM-based agents diagnose infeasibility in natural language [237], explain the details of decision making [238, 239], or interpret optimization processes and outcomes for end users [240, 241, 242].
Looking ahead, optimization can be reimagined as a multi-agent process, where LLMs assume complementary roles (e.g., problem formulation agents, solver orchestration agents, explanation agents, and evaluation agents) working together in dynamic teams. Such ecosystems would not only integrate soft knowledge that is difficult to formalize mathematically but also maintain transparency and interpretability for human collaborators. By evolving toward agentic ecosystems, the field can transcend the limitations of single-model pipelines, achieving adaptive, collaborative, and human-aligned optimization in complex real-world environments.
VIII Conclusion
This survey has presented a comprehensive and systematic review of the emerging field of LLMs for optimization. As an interdisciplinary synthesis of LLMs and optimization methods, this field is reshaping the foundations of optimization research, bringing new perspectives to long-standing challenges and inspiring a series of novel paradigms. By structuring the discussion from modeling to solving, we have aimed to clarify how LLMs can participate in, and ultimately transform, the entire optimization workflow.
On the modeling side, we examined how LLMs translate unstructured natural language into formal mathematical formulations, outlining the main workflows and highlighting the shift from prompt-based to learning-based methods. On the solving side, we proposed a three-level taxonomy grounded in the interaction between LLMs and evolutionary algorithms: (i) LLMs as stand-alone optimizers, (ii) low-level LLMs embedded into algorithmic components, and (iii) high-level LLMs serving as orchestrators for algorithm selection and generation. Beyond these methodological perspectives, we discussed how LLMs for optimization are being applied across computer science, the natural sciences, and engineering and industry. Representative applications such as neural architecture search, protein and molecular design, and wireless communication optimization demonstrate both the versatility of the field and its potential for real-world impact.
Looking forward, we envision LLMs for optimization evolving from isolated tools into dynamic, self-evolving, and agentic systems that can close the loop between modeling and solving, adapt continuously to real-time feedback, and collaborate with humans and other agents in complex environments. By outlining these trajectories, we hope this survey will serve not only as a systematic map of current progress but also as a catalyst for deeper exploration and innovation in this rapidly expanding domain.